### v1

# ACE: Abstractions for Communicating Efficiently

Abstract
A central but unresolved aspect of problem-solving in AI is the capability to introduce and use abstractions, something humans excel at. Work in cognitive science has demonstrated that humans tend towards higher levels of abstraction when engaged in collaborative task-oriented communication, enabling gradually shorter and more information-efficient utterances. Several computational methods have attempted to replicate this phenomenon, but all make unrealistic simplifying assumptions about how abstractions are introduced and learned. Our method, Abstractions for Communicating Efficiently (ACE), overcomes these limitations through a neuro-symbolic approach. On the symbolic side, we draw on work from library learning for proposing abstractions. We combine this with neural methods for communication and reinforcement learning, via a novel use of bandit algorithms for controlling the exploration and exploitation trade-off in introducing new abstractions. ACE exhibits similar tendencies to humans on a collaborative construction task from the cognitive science literature, where one agent (the architect) instructs the other (the builder) to reconstruct a scene of block-buildings. ACE results in the emergence of an efficient language as a by-product of collaborative communication. Beyond providing mechanistic insights into human communication, our work serves as a first step to providing conversational agents with the ability for human-like communicative abstractions.
1 Introduction
A key feature of human problem-solving is the use of abstractions. Humans use abstractions in many procedural tasks such as cooking, building shapes, and programming. In a seminal paper by Ho et al. (2019) that combines perspectives from Cognitive Science, AI and reinforcement learning, it is argued that abstractions guide learning, facilitate trade-offs, and simplify computation. They identified an outstanding open problem that “future work in both computer science and psychology will need to identify other pressures that can shape abstraction learning – such as the need to communicate and coordinate with others”. We carry out the first steps to solving this problem in the setting of co-operative artificial agents.
When presented with the opportunity for repeated interaction, human conversational partners tend towards more concise utterances (Hawkins, Frank, and Goodman 2020; McCarthy et al. 2021; Krauss and Weinheimer 1964). This tendency towards brevity is facilitated by the introduction of new, more informative abstractions into their shared language, which enables shorter utterances and thereby promotes more efficient communication and cooperation. How could AI agents be designed to introduce and exploit abstractions in a similar way in co-operative problem-solving?
Figure 1: In early rounds of the architect-builder game, the architect messages refer to horizontal or vertical blocks (a). After multiple interactions, the architect tries to introduce an abstraction (b), which after a learning period allows for shorter communication to solve the task (c).
We take a novel neuro-symbolic approach to study the introduction of abstractions in the context of emergent communication (EC) in multi-agent systems (Lazaridou and Baroni 2020; Foerster et al. 2016). We investigate this as a signalling game, a foundational tool in cognitive science (Lewis 1969; Regier, Kemp, and Kay 2015) to establish how human languages are optimised to communicate mental concepts efficiently. Our setting is an extension of the Architect-Builder game (McCarthy et al. 2021), depicted in Figure 1. This game features two agents who must complete a collaborative building task. The architect observes a goal-scene and instructs the builder to place horizontal and vertical blocks to recreate it. The architect must communicate this goal-scene using only linguistic utterances from a finite vocabulary. Initially, the architect’s messages refer to the placement of single horizontal or vertical blocks, as in Figure 1a. After a number of interactions, the architect can introduce new higher-level abstractions for common shapes, as in 1b, allowing for messages carrying more information, see 1c. For example, an utterance to stand for the “O shape” rather than having to refer to each individual block. Using such abstractions can significantly increase the efficiency of the collaboration as observed in human interactions.
Unlike traditional program synthesis which is limited to a fixed language, we draw on ideas from neuro-symbolic program synthesis and library learning (Ellis et al. 2021; Bowers et al. 2023) to introduce new abstractions into the language. These correspond to common sequences of instructions that would maximally reduce the length of future programs. However, it is not desirable to only naïvely consider compression, this would result in a huge language with a word for every single shape – even those rarely seen. The introduction of a new abstraction also incurs a cognitive cost in learning the modified language. For humans, there is typically a trade-off between competing pressures: to be informative and precise while not having to memorise too many obscure words causing a high cognitive load (Xu, Liu, and Regier 2020). To address this trade-off in our agents, we use techniques from emergent communication and reinforcement learning previously shown to promote these pressures in other settings (Carlsson, Dubhashi, and Johansson 2021; Carlsson, Dubhashi, and Regier 2023).
We test our model within an extended version of the Architect-Builder game comprising 961 goal-scenes and find that it exhibits the same tendency as humans — a development toward more informative and concise communication. Interestingly, we find that after a number of abstractions have been introduced, our model naturally converges to a stable language, situating this observation within the cognitive science literature on efficient communication (Kemp and Regier 2012; Gibson et al. 2017; Zaslavsky et al. 2019; Gibson et al. 2019). Our approach addresses many limitations of existing approaches by McCarthy et al. (2021) and Jergéus et al. (2022) and provides a valuable framework for future exploration in this area.
The contributions of our paper are as follows:
  • Responding to the challenge in (Ho et al. 2019), we initiate the study of abstractions in making communication between artificial agents engaged in a collaborative task more efficient.
  • We introduce Abstractions for Communicating Efficiently (ACE) which features a novel integration of techniques from EC, library learning and bandits.
  • We demonstrate how bandit techniques can be useful to control the exploration-exploitation trade-off when introducing new abstractions.
  • We provide an extended version of the architect-builder game initially proposed in McCarthy et al. (2021).
  • We demonstrate that ACE exhibits similar tendencies to humans and that the abstractions that it learns are subject to pressures for efficient communication.
  • We call for abstractions to be given greater attention within research on intelligent agents. Through doing so we anticipate opportunities for greater synergy in cooperative human-machine settings.
2 Architect-Builder Game
In the Architect-Builder Game, the architect is provided with a set of goal-scenes consisting of horizontal and vertical blocks on a grid as in Figure 1. The architect needs to communicate instructions to the builder (who does not know the goal state) that allow the builder to construct the goal-scene starting from a blank grid. This is inspired by the human experiments in McCarthy et al. (2021). In our work, the architect is a neuro-symbolic agent and the builder is a neural agent.
The architect is initialised with a pre-defined symbolic policy $\pi_{arch}$, which for each goal-scene contains a program to construct it. These programs consist of a sequence of instructions in a language ${\cal A}$, initially containing two primitive actions: placing either a horizontal or vertical block somewhere in the grid.
The agents are trained together to learn policies to enable the architect to communicate instructions to the builder to reach goal states from the initial empty grid while introducing useful abstractions, shortening future communications. We refer to the architect and builder’s communication policies as $\pi_{comm}$ and $\pi_{bldr}$, respectively.
We compose the goal-scene dataset $G$ similarly to McCarthy et al. (2021), where each scene is composed of two shapes placed side by side that are in turn built exclusively of the two primitive building blocks. These shapes are of different sizes and resemble either uppercase or lowercase letters from the English alphabet (see Appendix A.1), resulting in $31$ unique shapes, with multiple sub-shapes reoccurring in different shapes. The dataset results in a total of $961$ goal-scenes. This is over $100$ times bigger than that of McCarthy et al. (2021), which enables training neural agents. We split the dataset into 930 training scenes and 31 test scenes. These splits are constructed to ensure the distribution over shapes is the same. However, in the test set, the ordered pair of shapes constituting each goal-scene, does not appear in the training set.
3 Abstractions for Communicating Efficiently
There are three main components to ACE:
Abstraction
The architect identifies new abstractions: these are common sub-sequences within programs that are introduced through a library learning mechanism.
Communication
The (neural) communication between the architect and builder, using techniques from EC to learn a language for collaboration.
Exploration/Exploitation
Bandit techniques to choose between alternative programs for a given goal-scene.
3.1 Introducing Abstractions
The architect initially has a program consisting of primitive instructions to create a given goal-scene from an empty grid. As the interaction proceeds with the builder, the architect may discover new abstractions which can shorten the communication i.e. common sequences of instructions that recur during the interaction. This new abstraction can be used to rewrite the current program allowing for a shorter, compressed version which is then available as an alternative program for use in later interactions. Thus, at any stage in the interaction, the architect can choose among a set of available programs for each goal, each using a different combination of actions.
In order to find a more informative language the architect needs a way to introduce new terms into ${\cal A}$. The architect can analyse the available programs and introduce a new abstraction for a sub-sequence that enables the maximal reduction in the length of the whole set of programs. We choose a sub-routine according to a Bayesian procedure inspired by library learning-based neuro-symbolic architectures like DreamCoder (Ellis et al. 2021). We define this procedure below.
The architect evaluates a set of candidate sub-sequences and picks the one which maximises Equation 1:
$P(\mathcal{A}\cup\{a_{cand}\}|\{p_{i}\}_{i=1}^{N})\propto\prod_{i=1}^{N}P(p_{i}|\mathcal{A}\cup\{a_{cand}\})$ (1)
[…]
3.2 Communicating a program
Human languages are subject to pressures to be informative and to minimise cognitive load (i.e. to have a small language). In previous work, emergent communication methods have already been used to explain the formation and structure of human languages in domain such as numerals (Carlsson, Dubhashi, and Johansson 2021) and colours (Carlsson, Dubhashi, and Regier 2023). We apply EC techniques in our domain to study how pressures upon communication manifest within the context of abstraction learning.
Given a program, $p$, selected by the architect to communicate to the builder, a one-step signalling game is created for each instruction of $p$ which is of the form $(x,a,x^{\prime})$: grid state $x$ is transformed to grid state $x^{\prime}$ by the action $a$. We note that along with the action we pass positional information to the builder directly, we assume the encoding of this is pre-determined and we omit it from further notation. Each game involves an individual action being communicated from the architect to the builder which then attempts to infer the intended modification to the grid. In line with practices from EC, the architect’s and builder’s policies $\pi_{comm}$ and $\pi_{bldr}$ are both learnable and are parameterised by fully-connected neural networks. Model inference is achieved by the builder according to Equation 3, where, $\widehat{x^{\prime}}$ refers to an estimate of the next grid-state, $x^{\prime}$. Since our messages are discrete we use the gumbel-softmax relaxation to sample from discrete messages which makes our model end-to-end differentiable (Jang, Gu, and Poole 2017). Both the communication as well as builder models are trained to minimise the binary cross entropy loss given in Equation 4, where $U\times V$ denote the dimensionality of the grid. The accuracy of the reconstruction is used in the bandit’s reward function.
$\widehat{x^{\prime}}=\pi_{bldr}(x,\pi_{comm}(a))$ (3)
$\mathcal{L}(x^{\prime},\widehat{x^{\prime}})=-\frac{1}{UV}\sum_{i=1}^{|U|}\sum_{j=1}^{|V|}[x^{\prime}_{ij}\log(\widehat{x^{\prime}}_{ij})+\\ (1-x^{\prime}_{ij})\log(1-\widehat{x^{\prime}}_{ij})]$ (4)
3.3 Exploration/Exploitation of Abstractions
As abstractions are introduced, there will be multiple alternative programs of different lengths available to the architect. They may differ in reconstruction accuracy, in particular, programs containing a new abstraction will initially have low accuracy. If we want the agents to learn how to use these new abstractions they must be exposed to them. Hence, this is a classic exploration vs. exploitation problem which bandit techniques are well-suited for (Sutton and Barto 1998).
We view the selection between programs as a contextual bandit with combinatorial actions (Lattimore and Szepesvári 2020). We update the estimated value of a single action via Equation 5,
$Q(a)\leftarrow Q(a)+\alpha(r-Q(a))$ (5)
where $r$ is the reconstruction accuracy which is 1 if the action is successfully interpreted by the builder and 0 otherwise. The parameter $\alpha$ is the learning rate.
The value of an entire program is calculated according to Equation 6:
$Q(p)=\prod_{a\in|p|}\gamma Q(a)$ (6)
which captures the trade-off between program length and the communicative accuracy of instructions. The value $\gamma$ is in the semi-closed interval $(0,1]$ where a lower value results in a higher Q-value for shorter programs. Exploration vs. exploitation is undertaken using an $\epsilon$-greedy approach, where $\epsilon$ is held at a constant value, ensuring a constant level of exploration.
Through this, we allow the architect to control the regret, $r^{*}-r$, the difference between the optimal and observed reward, arising from the introduction of new abstractions. It is not obvious how the initialisation of the Q-values (Equation 5) may impact regret. Compelling cases can be made for both optimistic (a more aggressive introduction) and pessimistic (a more gradual introduction) strategies. Empirically (see Appendix A.4), we find the pessimistic strategy works better in our case.
3.4 Full Algorithm
Algorithm 1 shows how the components of ACE integrate. The algorithm alternates between the Communication Phase and the Abstraction Phase for $s$ steps, where $s$ is determined experimentally [^1] We found that 20 steps were a good choice for our setting, after which the language stabilises and no more abstractions are introduced in the emergent language..
The Communication Phase is repeated for $e$ iterations, determined experimentally. In each iteration, programs are sampled from $\pi_{arch}$ by the bandit’s $\epsilon$-greedy strategy and are used to form a dataset of grid-state transitions. This dataset forms the basis of our signalling game and is used to train $\pi_{comm}$, $\pi_{bldr}$ and to update the $Q$-values of actions. Here, we expect a language to emerge which is subject to pressures for efficient communication.
In the Abstraction Phase sub-sequences in the programs are identified. The sub-sequence that maximises Equation 1 is selected. This abstraction, $a_{new}$, is introduced into $\cal A$. $a_{new}$ may allow existing programs to be rewritten: in this case the resulting shorter programs are introduced into $\pi_{arch}$ for the respective goal-scene. Finally, $Q(a_{new})$ is initialised. More detailed descriptions of ACE’s phases can be found in Appendix A.2.
Algorithm 1 Abstractions for Communicating Efficiently (ACE)
0:  initial $\pi_{arch}$, initial $\cal A$, $q_{init}$, $\alpha$, $\epsilon$, $e$, $s$
0:  trained $\pi_{comm}$, $\pi_{bldr}$, $Q$, updated $\pi_{arch}$, $\cal A$
1:  Initialise neural networks $\pi_{comm}$ and $\pi_{bldr}$
2:  Initialise action Q-values $Q(a)$ to $q_{init}$, $\forall a\in\mathcal{A}$
3:  for $s$ steps do
4:     // Communication Phase
5:     for $e$ rounds do
6:        For each goal choose a program in $\pi_{arch}$ via bandit
7:        Create a training dataset $D$ of triples $(x,a,x^{\prime})$ of transitions based on chosen programs
8:        Play signalling games with $D$ to train $\pi_{comm}$, $\pi_{bldr}$ and $Q$-values using Eq. 3, 4 and 5
9:     end for
10:     Optional: evaluate on hold-out test dataset
11:     // Abstraction Phase
12:     Find action sub-sequences in programs
13:     Choose new abstraction $a_{new}$ that maximises Eq. 1
14:     Compress existing programs using $a_{new}$, if possible
15:     For each goal, add new programs to $\pi_{arch}$
16:     Add $a_{new}$ to $\cal A$ and initialise $Q(a_{new})$ to $q_{init}$
17:  end for
4 Empirical Evaluation
To evaluate ACE we investigate the following questions: 1) Does ACE result in a final language allowing for shorter communications? 2) Does communicative pressures impact what abstractions get adopted?
The behaviour of ACE is compared to two naïve baselines. The first, No abstractions, is ACE without the abstraction phase. The second, Greedy, always picks the shortest program. We use reward (also referred to as reconstruction accuracy) and average program length to compare these approaches. Unless otherwise stated results obtained are averaged over $16$ runs. We provide further implementation details in Appendix A.3.
Figure 2: Average program length vs. steps for ACE, Greedy and No abstractions. Line indicates mean performance and shaded regions indicate the 95% confidence interval.
Figure 3: Test reward vs. steps for ACE, Greedy and No abstractions. Line indicates mean performance and shaded regions indicate the 95% confidence interval.
4.1 ACE reduces program length
ACE does indeed converge to shorter programs over interaction steps, as seen in Figure 2. When we compare ACE to the No abstraction variant, the program lengths are halved, going from an average length of $10$ to one of $4.92\pm 0.20$, indicating that a more efficient language has been derived. The Greedy variant reduces program lengths even more, however, this comes at a price. In Figure 3 we see drastic drops in accuracy for Greedy every time an abstraction is introduced, whereas ACE’s slower introduction leads to smaller drops in accuracy. For a while, both strategies manage to recover near full accuracy, but after some time, the Greedy communication fails to recover and becomes progressively worse.
4.2 The Communicative Pressures on Abstractions
Figure 4: The abstractions introduced by ACE in an example run. We circle in green the abstractions that are adopted by the architect in its final language.
(a) The first abstraction, resembling a tower, being introduced successfully.
(b) The thirteenth abstraction, resembling an X, is introduced, but not retained.
Figure 5: Different abstractions being introduced. We show the mean and 95% confidence interval for Q-value (black) and frequency (blue) versus epochs (of which there are $40$ in a step) over $20$ random seeds.
Figure 4 shows the abstractions that ACE introduces in an example run [^2] Further examples, which are quite similar, can be found in our code repository.. Abstractions circled in green are in the final language. ACE is able to compose abstractions: smaller abstractions may be utilised later in larger abstractions. For example, abstraction 17 is a sequence of two instances of abstraction 3.
We note that not all of the 20 proposed abstractions are retained in the final language. In Figure 5(a) and 5(b), we show abstractions that experience different fates. Figure 5(a), shows the Q-value and frequency of abstraction 1 from Figure 4 (a tower), which was retained. The tower is very common in our dataset, appearing $1380$ times. Figure 5(b), shows the same for abstraction 13 from Figure 4 (the letter X), which was not retained. The X is relatively uncommon and appears only 60 times. In the case of the tower, after a period of exploration, ACE readily adopts it. We note a brief period of instability that occurs at step $2$, while a second abstraction is being introduced, after which the usage of the tower gradually stabilises. The X on the other hand, has a small spike in frequency initially, but it quickly drops to zero as the agents abandon it, having found it too hard to learn.
In Figure 6, we show how the composition of our language changes throughout the repeated interactions. The relative proportion of primitives is reduced in favour of abstractions referring to either one of the 31 discrete shapes in the dataset (roughly 40%), or a sub-shape (roughly 20%). We show how this change in distribution impacts the composition of programs in Figure 7 with two goal-scenes examples. We see that the introduction of new abstractions enables for programs to be rewritten much more compactly.
Figure 6: Relative proportion of actions as the language changes over training. Primitives refer to the initial actions, shapes refer to one of the 31 discrete shapes appearing in goal-scenes, and sub-shapes refer to anything else.
Figure 7: Two goal scenes with representations of their initial and final programs chosen by ACE.
5 Discussion
Addressing the challenge in (Ho et al. 2019), we initiated the study of how abstractions could be used to make communication between artificial agents more efficient. In contrast to (McCarthy et al. 2021) who studied this question with human subjects, ours is the first to explore it in the setting of artificial agents.
We have demonstrated that ACE exhibits the capability of developing a compact collaborative language, resulting in shorter programs. The size of the language is the consequence of pressures that naturally arise through the need to communicate about a shared task. This is achieved by ACE through a novel combination of EC, library learning and bandits. Our work serves as a bridge between abstraction learning and efficient communication (Kemp and Regier 2012; Gibson et al. 2017; Zaslavsky et al. 2019; Gibson et al. 2019; Denić and Szymanik 2024). It builds on previous research showing that Emergent Communication can be used to learn efficient communication in specific domains, such as numerals (Carlsson, Dubhashi, and Johansson 2021) and colors (Carlsson, Dubhashi, and Regier 2023), and extends this to the learning of abstractions.
Beyond importance within cognitive science, we believe our results have real implications for machine learning. Significant research effort is being invested into exploring the role of natural language as a mechanism for humans to provide instructions to intelligent agents (Brohan et al. 2023; Shi et al. 2024). It is natural for humans to introduce abstractions and we should create systems that can accommodate this. Research promoting this capability can facilitate improved cooperation between humans and agents. We believe that ACE represents a step towards equipping intelligent agents with flexible and extendable languages.
Limitations
So far, we have not applied ACE to domains beyond the Architect-Builder game, and extension towards other settings, such as the ARC challenge (Chollet 2019) or PizzaCommonSense (Diallo et al. 2024) is left to future work. In novel environments such as these, our assumption that a set of initial programs exists for each goal-state may not be realisable. This would require also initial programs to be learned which would further complicate the task of learning to communicate. Furthermore, our work does not yet analyse ACE’s languages from an information-theoretic perspective, e.g. as a trade-off between average program length and vocabulary size, situating them in relation to a Pareto frontier.
Future Work
ACE presents a step towards understanding how abstractions are introduced for instruction sequences. It has been shown by Krauss and Weinheimer (1964) that abstractions also appear within repeated exchanges expressed in natural language. This is a complementary perspective on the role of abstractions within human communication. We intend to integrate both of these perspectives into a unified model using LLM-based agents like those in (Brohan et al. 2023).
6 Related Literature
Emergent communication in the Architect-Builder game was first studied in humans by McCarthy et al. (2021), however only on a smaller dataset with only 12 goal-scenes. They also propose a simple Bayesian model for the task, inspired by Goodman and Frank (2016) and Ellis et al. (2021). However, this is not a proper multi-agent system and is limited to assuming that abstractions are introduced automatically by a centralised process in both the architect and builder at the same time. Our work addresses this major limitation by situating the task in a multi-agent system where the introduction of abstractions is controlled by the architect, a more plausible representation. When a new abstraction is introduced, ACE allows the architect and builder to negotiate a change to their common language to incorporate it, or if it proves difficult to learn, discard it. Furthermore, our setting does not require an explicit parameter to keep the language size manageable, this appears automatically as a consequence of communicative pressures in ACE. Jergéus et al. (2022) conduct a preliminary study into using multi-agent reinforcement learning for the Architect-Builder task, but make many simplifications. Most notably, there is no learning of a messaging policy and the abstractions are assumed to be immediately understood by the builder. Under this assumption, this actually collapses to a single agent setting and is not subject to pressures for efficient communication.
While EC has been used to study areas such as the impact of populations (Chaabouni et al. 2022; Kim and Oh 2021; Rita et al. 2022), compositionality (Mordatch and Abbeel 2018; Chaabouni et al. 2020; Ren et al. 2020), and naming conventions in semantic categories (Carlsson, Dubhashi, and Regier 2023; Carlsson, Dubhashi, and Johansson 2021), we introduce it as a tool to study the process of abstraction and combine it with library learning and bandits.
7 Conclusions
When human conversational partners repeatedly cooperate on a task, they tend to develop more concise language. In this paper, we study this process within a collaborative building game known as the architect-builder game. We propose a mechanistic explanation that integrates bandit models with combinatorial actions, emergent communication, and library learning to model the abstraction process. Together, these components enable for the proposal and evaluation of new abstractions based on their communicative efficiency and their impact on program length. We find that our method aligns with human tendencies towards more concise language over repeated interactions. In comparison to previous approaches, our model represents a more naturalistic approach, which removes restrictions on the manner in which abstractions are introduced and naturally converges to a fixed policy as a result of emergent communication thereby reducing dependence on hard-coded parameters.
Appendix A Appendices
A.1 Dataset
Figure 8: The $31$ shapes that in pair compose each goal-scene.
Each goal-scene in the dataset is composed of two shapes placed next to each other, as in Figure 1. These are sampled from 31 shapes we design and report in Figure 8. Each shape is meant to resemble an uppercase or lowercase letter from the English alphabet. We note that there are multiple subshapes that recur several times within the 31 shapes, making our dataset a prime candidate for studying the introduction of abstraction.
A.2 Algorithms composing ACE
In this Section, we provide more detailed algorithms for the focal points of ACE’s methodology. Algorithms 2 and 3 refer to the Communication Phase, which we break down further in two parts: the dataset generation for the signalling game and the communication round done via EC. Algorithm 4 instead refers to the Abstraction Phase.
Algorithm 2 Dataset generation via bandit
0:  $\pi_{arch}$, $Q$, bandit strategy
0:  $D$ A dataset of signalling games
1:  $D=\{\}$
2:  for $g\in G$ do
3:     Get set of programs for $g$ from $\pi_{arch}$
4:     Compute Q-values for programs (Equation 6)
5:     Pick program, $p$, via bandit strategy
6:     break $p$ in a set of triples $(x,a,x^{\prime})$ of transitions
7:     $D\leftarrow D\bigcup\{(x_{i},a_{i},x^{\prime}_{i})\}_{i\in|p|}$
8:  end for
Algorithm 3 A round of communication of ACE
0:  $D$, $\pi_{comm}$, $\pi_{bldr}$, $Q$
0:  trained $\pi_{comm}$, $\pi_{bldr}$, $Q$
1:  # Play signalling game and update policies
2:  while $D$ is not empty do
3:     Sample batch $B$ from $D$ without replacement
4:     $loss=0$
5:     for $(x,a,x^{\prime})\in B$ do
6:        $m=\pi_{comm}(a)$
7:        $\widehat{x^{\prime}}=\pi_{bldr}(x,m)$
8:        $loss\leftarrow loss+\mathcal{L}(x^{\prime},\widehat{x^{\prime}})$
9:        Update $Q(a)$ for all $a$ in $B$ via Equation 5
10:     end for
11:     Update $\pi_{comm}$ and $\pi_{bldr}$ via gradient descent
12:  end while
Algorithm 4 Abstraction phase
0:  $\pi_{arch}$, $\cal A$
0:  updated $\cal A$ and $\pi_{arch}$
1:  $Z=\{\}$ The set of preferred programs
2:  for $g\in G$ do
3:     Get set of programs for $g$ from $\pi_{arch}$
4:     Compute Q-values for programs (Equation 6)
5:     Pick programs, $p$, via greedy bandit strategy
6:     $Z\leftarrow Z\bigcup\{p\}$
7:  end for
8:  Find all sub-sequences in $Z$ (candidate abstractions)
9:  Select best abstraction, $a_{new}$, via Equation 1
10:  $A\leftarrow A\bigcup\{a_{new}\}$
11:  for $g\in G$ do
12:     for $p\in\pi_{arch}(g)$ do
13:        if $a_{new}$ is applicable to $p$ then
14:           Add $p$ re-written in terms of $a_{new}$ to $\pi_{arch}(g)$
15:        end if
16:     end for
17:  end for
A.3 Implementation Details
We implement ACE within Python version 3.9.18 and PyTorch version 2.2.1 (Paszke et al. 2019) and run experiments on Apple M2 Pro CPU (16 GB) on MacOS 14.4.1. The communication policies, $\pi_{comm}$ and $\pi_{bldr}$, are deep neural networks which have $1$ and $2$ hidden layers, respectively. Each hidden layer contains $200$ neurons and utilises ReLU activation functions. The outputed layer of $\pi_{comm}$ is comprised of $30$ neurons, which is its maximum vocabulary size. The output layer of $\pi_{bldr}$ is comprised of $81$ neurons, which is the size of the grid. Each output neuron in $\pi_{bldr}$ represents the builder’s belief that the grid index is occupied. We use the gumbel-softmax relaxation with the hard parameter so that the networks can be differentiated end-to-end. The model is optimised to minimise binary cross entropy error and we use a learning rate of $0.0009$ in conjunction with the ADAM optimiser (Kingma and Ba 2014). The bandit hyperparameters $\alpha$, $\gamma$, $\epsilon$ and $q_{init}$ are set to $0.5$, $0.99$, $0.1$, and $0.0$ respectively. Note, that before an abstraction is introduced we prune all but the $3$ best programs (determined empirically) for each goal-scene in $\pi_{arch}$ to not incur an exponential growth in the number of programs for each goal-scene. We also set the number of epochs $e$ to 40. All hyperparameters chosen are experimentally determined through grid search.
A.4 Q-value initialisation: Pessimism is best
In §3.2, we considered the impact of initialisation on learning within ACE. Here, we establish its impact on learning of the common “tower” abstraction. We compare a pessimistic strategy ($q_{init}=0$) to an optimistic strategy ($q_{init}=1$) in terms of cumulative regret, defined as $Regret_{T}=\sum_{t=1}^{T}\left(r^{*}-r_{t}\right)$. In Figure 9, we see that a pessimistic initialisation of 00 incurs significantly less cumulative regret than an optimistic value of $1$. This is attributable to the optimistic variant’s tendency to force the abstraction into use before the builder has had the opportunity to become familiar with it. The pessimistic agent introduces the new abstraction more gradually while still providing the opportunity for learning of a common language.
Figure 9: Cumulative regret for an optimistic and pessimistic initialisation of Q-values described in Equation 5. Line indicates mean performance and shaded regions indicate the 95% confidence interval for $16$ random seeds. Epoch interval is between step $1$ and step $2$.
Acknowledgments
This work was supported by funding from from the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation and the Swedish Research Council. The computations in this work were enabled by resources provided by the Swedish National Infrastructure for Computing (SNIC). We also extend our thanks to Emil Carlsson for insightful discussions at various points throughout the project.